the ordered pair (1, -2) is a solution of the following system..true or false?
x+y=-3
x-y=1

Answer by checkley71(8403)

1-2=-3
-1 DOES NOT=-3
1-(-2)=1
1+2=1
3 DOES NOT=1
THE ANSWER IS FALSE.